Can someone help me with these 2, I am having trouble solving them.

Question 4

Using the tables for water determine the specified property data at the indicated state.
For H2O at p = 0.7 MPa and v = 0.62 m3/kg, find h in kJ/kg.
Note: Round up the answer to 2 decimal places.

---------------------------------------------------------------------

Question 5

Using the tables for water determine the specified property data at the indicated state.
For H2O at p = 0.7 MPa and v = 0.57 m3/kg, find u in kJ/kg.
Note: Round up the answer to 2 decimal places.
Answer #1

From Steam tables at 0.6MPa ,at v = 0.62m3/kg, h = 3579kJ/kg

at 0.8MPa and at v = 0.62m3/kg h = 4156kJ/kg

at 0.7MPa , by linear interpolation, h= (3579+4156)KJ/kg = 3876.50KJ/kg

From Steam tables at 0.6MPa ,at v = 0.57m3/kg, h = 3415kJ/kg

at 0.8MPa and at v = 0.62m3/kg h = 3962kJ/kg

at 0.7MPa , by linear interpolation, h= (3579+4156)KJ/kg = 3688.50KJ/kg

The steam table values may change little bit from table to table, use the values given in your table to obtain the required answer.